# Modular reduction of complex representations of finite reductive groups

### v1

@@ 8. Computations with the elements $M_{w}$
In [L, Section 2], Lusztig computed by hand the elements $M_{w}$ in Types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, and $A_{3}$. The $M_{w}$ which we define in the present paper agree with those appearing in loc. cit. We reproduce these computations below.
Our definition gives an explicit algorithm for computing $M_{w}$ in any type. Using SageMath, we implemented this algorithm and obtained explicit expansions for $M_{w}$ in the basis of Weyl characters in Types $B_{3},A_{4},B_{4},D_{4}$. In Tables 3 and 5, we give such expressions in Types $B_{3}$ and $A_{4}$ to provide some additional examples.
Table 1. A table of $M_{w}$ for Types $A_{1},A_{2},B_{2},G_{2}$, previously computed in [L, Section 2].
@@
Table 2. A table of $M_{w}$ in Type $A_{3}$, with each $M_{w}$ being written first in the basis of Weyl modules $V_{\lambda}$ and then in the basis of irreducibles $L_{\lambda}$. The expansion in terms of $V_{\lambda}$ also appears in [L, Section 2].

### v2

@@ 8. Computations with the elements $M_{w}$
In [undefp, Section 2], Lusztig computed by hand the elements $M_{w}$ in Types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, and $A_{3}$. The $M_{w}$ which we define in the present paper agree with those appearing in loc. cit. We reproduce these computations below.
Our definition gives an explicit algorithm for computing $M_{w}$ in any type. Using our Sage implementation [undefu], we obtained explicit expansions for $M_{w}$ in the basis of Weyl characters in Types $B_{3},C_{3},A_{4},B_{4},D_{4}$. The repository can be used to reproduce all of the computations in the present paper and to work with the constructions introduced here. In Tables 3, 4, and 5, we give such expressions in Types $B_{3},C_{3}$ and $A_{4}$ to provide some additional examples.
Table 1. A table of $M_{w}$ for Types $A_{1},A_{2},B_{2},G_{2}$, previously computed in [undefp, Section 2].
@@ 11 $V_{0,0}$
Table 2. A table of $M_{w}$ in Type $A_{3}$, with each $M_{w}$ being written first in the basis of Weyl modules $V_{\lambda}$ and then in the basis of irreducibles $L_{\lambda}$. The expansion in terms of $V_{\lambda}$ also appears in [undefp, Section 2].